PRESENT VALUE Word Problems

Note: The Present Value Formula is PV= FV/ (1+r)^n


1. Angelo wants to renovate his house in 3 years. He estimates the cost to be P300,000. How much must Angelo invest now at 8% interest compounded quarterly, in order to have P300,000, 3 years from now.

2. Arnold and Azura want to save P500,000 in 5 1⁄2 years to renovate their rest house. If the bank of Amber is paying 8% interest compounded quarterly, how much must they deposit now in order to have the money for the project?

3. Find the present value of P85,000, if the interest rate is 6% compounded quarterly, for 10 years.

4. Ashley and Aileen want to accumulate P300,000, 17 years from now, as a college fund for their baby daughter, Aria. Use the present value formula to calculate how much they must invest now, at an interest rate of 8% compounded semi-annually, in order to have P300,000 in 17

years.


1
Expert's answer
2021-03-31T13:03:14-0400


1.

PV= FV/ (1+r/n)nt

PV= P300000/(1+0.08/4)12

=P236,547.95


2.

PV= FV/ (1+r/n)nt

PV= P500000/(1+0.08/4)22

=P323,419.52


3.

PV= FV/ (1+r/n)nt

PV= P85000/(1+0.06/4)40

=P46857.30


4.

PV= FV/ (1+r/n)nt

PV= P300000/(1+0.08/2)34

=P79065.63